A weightlifter works out at the gym each day. Part of her routine is to lie on her back and lift a 45 kg barbell straight up from chest height to full arm extension, a distance of 0.52 m .

How much work does the weightlifter do to lift the barbell one time?

If the weightlifter does 23 repetitions a day, what total energy does she expend on lifting, assuming a typical efficiency for energy use by the body (25%).

How many 500 Calorie donuts can she eat a day to supply that energy?
Answer #1

work done during one lift=mgh=45*9.81*0.52=229.32 J  

work done in 23 lifts= 23*229.32=5.274KJ

1 cal =4.18 J so work done= 5274.36 J=1261.8 cal

also efficiency =25% so energy required=1261.8/0.25=5047.2 cal